A Riverboat Cruise Line boat travels 8 mi upstream and 8 mi downstream. The total time for both parts of the trip is 3 hr. The speed of the stream is 2 mph. What is the speed of the boat in still water?
Answer by mananth(16949)

let boat in still water be x mph
stream rate = 2mph
...
upstream rate = x-2
downstream rate = x+2
...
8/(x-2) +8/(x+2) =3
8(x+2)+8(x-2) / x^2-4 =3
8x+16+8x-16 = 3x^2-12
16x=3x^2-12
3x^2-16x-12=0
3x^2-18x+2x-12=0
3x(x-6)+2(x-6)=0
(x-6)(3x+2)=0
x= 6 speed of boat in still water
...
m.ananth@hotmail.ca